OpenStudy (anonymous):

square root of 6 times the square root of 3 divided by the square root of 9

OpenStudy (anonymous):

\[\sqrt{2}\] i believe

OpenStudy (anonymous):

root of 6*root of 3 = root of 18 this root of 18 is written as 3 root 2 root of 9 = 3 then the 3 cancel out and the final answer is sqaure root of 2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

U may combine sqrt(6*3/9) = sqrt(18/9) = sqrt(2).

OpenStudy (anonymous):

it's about difference two squares use difference of two squares

OpenStudy (anonymous):

@agradecompetent That is rubbish.